Two functions f and h are defined on the set of real numbers, R, by f(x) = (4−x^2)^1/2 and h(x) = (3−x)^1/ 2 (a) Find the domains of f and h. (b) Find foh(x).
1
Expert's answer
2020-07-12T15:49:02-0400

Part (a)

f(x)=(4-x^2)^(1/2)

Find the domain of f.

f(x)=(4-x^2)^(1/2)

4-x^2≥0

4≥x^2

-2≤x≤2

Domain of f: [-2, 2].

 

h(x)=(3-x)^(1/2)

Find the domain of h.

h(x)=(3-x)^(1/2)

3-x≥0

x≤3

domain of h: (-∞, 3].

 

Part (b)

Find foh(x)

foh(x)=f[h(x)]

and,

f(x)=(4-x^2)^(1/2), h(x)=(3-x)^(1/2)

 

f(h(x))={4-[(3-x)^(1/2)]^2}^(1/2)

f[(3-x)^(1/2)]=(1+x)^(1/2)

Therefore,

foh(x)=(x+1)^(1/2)